# Agent-X: Evaluating Deep Multimodal Reasoning in Vision-Centric Agentic Tasks

### v1

@@ Abstract
Deep reasoning is fundamental for solving complex tasks, especially in vision-centric scenarios that demand sequential, multimodal understanding. However, existing benchmarks typically evaluate agents with fully synthetic, single-turn queries, limited visual modalities, and lack a framework to assess reasoning quality over multiple steps as required in real-world settings. To address this, we introduce Agent-X, a large-scale benchmark for evaluating vision-centric agents’ multi-step and deep reasoning capabilities in real-world, multimodal settings. Agent-X features 828 agentic tasks with authentic visual contexts, including images, multi-image comparisons, videos, and instructional text. These tasks span six major agentic environments: general visual reasoning, web browsing, security and surveillance, autonomous driving, sports, and math reasoning. Our benchmark requires agents to integrate tool use with explicit, stepwise decision-making in these diverse settings. In addition, we propose a fine-grained, step-level evaluation framework that assesses the correctness and logical coherence of each reasoning step and the effectiveness of tool usage throughout the task. Our results reveal that even the best-performing models, including GPT, Gemini, and Qwen families, struggle to solve multi-step vision tasks, achieving less than 50% full-chain success. These findings highlight key bottlenecks in current LMM reasoning and tool-use capabilities and identify future research directions in vision-centric agentic reasoning models. Our data [^1] https://huggingface.co/datasets/Tajamul21/Agent-X and code [^2] https://github.com/mbzuai-oryx/Agent-X is publicly available.
@@ 1 Introduction
Agentic frameworks enable AI systems to interact with their environment by perceiving inputs, invoking tools, and executing actions. While perception and tool use are core components, solving complex tasks effectively requires reasoning, i.e., the ability to draw logical inferences, make decisions, and adapt over time based on multimodal inputs such as text, images, video, and temporal context [27, 70, 72, 60]. Deep reasoning in agents helps effectively plan, execute, and adapt across evolving scenarios [33, 43]. Recent works have explored integrating reasoning and tool use within large multimodal models (LMMs)[2, 6, 9, 55, 22, 17], where LMMs act as controllers for planning, and callable tools handle external actions[11, 18, 44]. This architecture enables agents to combine perception, visual understanding, symbolic reasoning, and generative capabilities, significantly improving task performance in complex environments [62, 38, 43].
@@
Existing benchmarks for agentic systems have primarily focused on text-based interactions, with limited support for multimodal inputs such as images, videos, and multi-image comparisons [38, 33, 62]. While some efforts have extended to multimodal tasks, they are often restricted to static images, synthetic environments, or narrowly scoped domains, offering limited insight into an agent’s ability to perform complex, tool-driven reasoning. A key limitation is inadequate evaluation of deep reasoning [^3] In our context, Deep reasoning refers to coherent, multi-step inference grounded in multimodal context.: current benchmarks either neglect this aspect entirely or lack principled metrics for assessing multi-step logical coherence [62, 43, 33, 59]. Furthermore, most existing benchmarks rely on either fully synthetic tasks or manual annotations, which lack scalability and fail to capture the depth of reasoning or the complexity of video-based multimodal interactions, falling short of modeling real-world agentic tasks [19, 60], as summarized in Table 1. As agents rapidly evolve in their capabilities, existing benchmarks struggle to keep pace, underscoring the need for more comprehensive, reasoning-centric evaluations to track planning, adaptation, and tool use in authentic, multimodal interactive scenarios [26].
@@
Table 1: Comparison of Vision-Centric Agent Benchmarks. Columns show key dimensions including scale, realism, modality, reasoning depth, tool interaction, and annotation quality. Our benchmark Agent-X uniquely supports all criteria with 828 diverse, manually verified agentic tasks.
@@
Benchmark Agentic Tasks # Tools Real-world Queries Multimodal Inputs Deep Reasoning Executable Tools Hybrid Annotation
APIBench [46] 3
APIBank [29] 73
ToolBench [48] 3451
MLAgentBench [20] - 8
GAIA [38] 466 26
SWE-Bench [23] -
GTA [62] 229 14
MLE-Bench [10] -
m&m’s [34] - 33
RE-Bench [63] 7 -
ScienceAgent [13] 102 -
MLGym [43] 13 16
Agent-X (Ours) 828 14
@@
To address these limitations, we introduce Agent-X, a large-scale evaluation benchmark for vision-centric agents that emphasizes two core principles: multimodal reasoning and vision-first evaluation. Agent-X rigorously tests agents’ ability to process complex visual and textual inputs, execute tool-augmented plans, and perform deep reasoning across real-world tasks. A primary feature of our benchmark is its deep reasoning assessment, i.e., judging the ability to perform coherent, multi-step problem-solving, logical inference, and adaptive planning. To move beyond surface-level evaluation (focused on the correctness of the final result), we introduce fine-grained metrics that capture both intermediate reasoning steps and overall task coherence. This approach helps distinguish genuine logical progression from inconsistencies and confabulations (plausible-sounding but disconnected steps) in reasoning chains, as illustrated in Figure 1. Agent-X spans six diverse multimodal environments, incorporating images, videos, and spatiotemporal contexts to evaluate agents in rich, realistic settings that demand generalization beyond text. Our task pipelines are derived from authentic user queries without explicit step instructions, simulating naturalistic agent interaction as highlighted in Table 2. These tasks are paired with executable toolchains via a semi-automated pipeline that integrates a broad set of real-world tools, enabling scalable tool-augmented decision-making. In Agent-X, these components create a comprehensive framework for evaluating agentic performance in perception, reasoning, and action. Our main contributions are as follows:
@@
ToolBench [48] I’m writing a blog post… You can first retrieve available figlet styles and then generate ASCII art for the strings using the style. figlet, list figlet styles
APIBench [46] I am an engineer at Uber and… Write a python program in 1 to 2 lines to call API in TorchHub. ObjectDetection
GTA[62] How much should I pay for the soda in the picture according to the price on the menu? ImageDescription, CountGivenObject, OCR
m&m’s [34] I need an illustration for my children’s book. I’ve imagined a scene where there’s a large group of little kids… After we have the image, we also need to identify all the objects, then add labels to them. ImageGeneration, ObjectDetection, Tagging
@@ 2 Related Work
Large Multimodal Agents: The rapid progress in LMMs [2, 17, 55, 9, 22, 6] has catalyzed the development of agentic frameworks capable of autonomous planning, tool use, and decision-making. A growing body of research has explored the extension of LMMs beyond conventional text generation by integrating them with external tools such as APIs [76], document processing [39], operating systems [37], and web interfaces [52], allowing them to interact meaningfully with their environments. This evolution has led to the emergence of tool-augmented agents such as Avatar [66], LangChain [11], AutoGPT [18], and BabyAGI [40], which provide frameworks for textual reasoning and execution of external actions. In the context of web browsing, systems such as WebShop [71], WebGPT [41], and WebCPM [47] enhance LMMs with capabilities for browsing, searching, and information retrieval. Beyond textual modalities, integrations such as RestGPT [51] and AppAgent [77] enable interaction with REST APIs and emulate touchscreen operations, respectively. Data-centric systems like DataCopilot [79] autonomously execute complex workflows. In contrast, agents such as HuggingGPT [50] and ModelScopeAgent [28] coordinate tasks across multiple specialized models. In the multimodal domains, vision-centric agents like MLLMTool [61] and LLaVA-Plus [32] equip LMMs with the ability to reason about visual content through pre-trained vision models. Nevertheless, there is no standardized evaluation protocol to rigorously assess the reasoning capabilities of these agents, particularly in tasks that require tool invocation and decision-making.
@@
Multimodal Agentic Benchmarks: A broad range of recent benchmarks have emerged to evaluate the performance of agentic frameworks. Recent efforts like ToolBench [48] introduce REST APIs and evaluate tool use with metrics such as pass rate and win rate, while APIBench [46] assesses accuracy for APIs. Complementing these, API-Bank [29] presents a diverse suite of commonly used APIs, such as search engines and hotel reservations, offering a comprehensive framework for evaluating the planning, retrieval, and execution skills of agents based on LMM. Benchmarks such as ToolQA [82], Gentopia [68], Gorilla [46], and AgentBench [33] often repurpose standard data sets or follow restricted evaluation protocols, which limit their utility in assessing open-ended and generalizable tool use capabilities. To address this, GAIA [38] introduces simple but conceptually challenging questions for human-like understanding, emphasizing general-purpose agentic intelligence. Similarly, GTA [62] evaluates agents with executable toolchains and real-world deployed tools in unconstrained environments, while Osworld [67] offers multi-step tasks inspired by actual user behavior in computer operating systems. The MLGym [43] benchmark further enhances realism by integrating complex multimodal tasks that require deep reasoning and tool use, simulating dynamic, real-world interactions. As shown in Table 1, despite promising advances, current evaluations face two key limitations: (1) most benchmarks focus on final answer accuracy, overlooking interpretability, which is crucial for agentic tasks involving reasoning and tool use, and (2) the absence of standardized and interpretable evaluation protocols hinders meaningful comparison across methods.
@@
Reasoning in Large Multimodal Models: Numerous methods have been introduced to enhance the reasoning capabilities of LMMs, enabling models to handle complex, multi-step problems through explicit reasoning processes [5, 69, 78, 60]. Following the introduction of reasoning techniques, various benchmarks have been developed in parallel to assess the quality and effectiveness of reasoning and chain-of-thought processes in models. Early benchmarks like CLEVR [24] focused on basic visual reasoning using synthetic scenes testing logic, counting, and comparisons. As tasks became more complex, StrategyQA [16] introduced multi-hop reasoning, while ScienceQA [49] incorporated multimodal input for scientific understanding. MathVista [36] unified 28 math-related datasets for visual math reasoning, and ShareGPT-4o [78] provided 193k chain of thought samples to strengthen intermediate reasoning in LMMs. The m&m’s [34] benchmark focuses on multi-step, multimodal reasoning and evaluates various planning strategies. However, many benchmarks, including m&m’s, rely on AI-generated queries with predefined tool sequences, limiting their realism. LlamaV-01 [60] introduced a benchmark focused on assessing tasks requiring multiple reasoning steps. While these reasoning-focused benchmarks have been introduced, they do not address the agentic framework or evaluate tool usage and the associated reasoning within vision-centric agentic tasks.
@@ 3 Agent-X Benchmark
The Agent-X benchmark rigorously evaluates the capacity of vision-centric agents to perform complex reasoning tasks that necessitate proficient tool usage across diverse, real-world scenarios. It accommodates a broad spectrum of input modalities, including single-image analysis, multi-image comparisons, video interpretation, and multimodal image-text interactions. Agent-X presents agents with authentic visual data and requires them to use multi-step reasoning processes. The construction of Agent-X follows a semi-automated pipeline (illustrated in Figure 2), where an initial set of candidate queries is generated through large multimodal models (LMM) and subsequently refined and validated by human experts.
@@ 3.1 Benchmark Design
In our framework, $\mathcal{T}_{c}$ comprises a collection of 14 tools distributed across diverse environments. A comprehensive list of these tools is presented in Appendix §B. The queries $\mathcal{Q}$ are classified into three distinct categories: factual, interpretive, and generative. Illustrative examples of these query types are provided in Appendix §H. In the case of a factual query, the answer is a specific, uniquely determined value, such as a number or a phrase. For an interpretive query, the final answer typically consists of descriptive text. While the answer is not unique, it conveys a general concept or idea, with the final response comprising a reference answer. For a generative query, we do not directly utilize the generated output since the LMMs are descriptive models; they only provide textual descriptions rather than generating visual content like images/videos. In these instances, $\mathcal{A}_{i}=\emptyset$.
@@ 3.2 Task Pipeline
Our query construction process begins with generating candidate queries using an LMM, which is provided with the visual input $\mathcal{V}_{i}$ and the available complete toolset $\mathcal{T}_{c}$. The LMM generates initial queries that human annotators then refine and validate for clarity, coherence, and realism. In the next step, LMM is re-engaged with the refined query to produce a detailed reasoning trace $\mathcal{R}_{i}$, capturing the sequence of tool calls (input/output), intermediate decisions, and final answer $\mathcal{A}$. Human reviewers further ensure that the trace is logically consistent and aligned with the query, while the final answer, $\mathcal{A}_{i}$ and its justification, $\mathcal{J}_{i}$ are validated for correctness and comprehensibility. The prompts provided to the LMM for query generation and reasoning trace construction are included in Appendix Appendix §I.

### v2

@@ Abstract
Deep reasoning is fundamental for solving complex tasks, especially in vision-centric scenarios that demand sequential, multimodal understanding. However, existing benchmarks typically evaluate agents with fully synthetic, single-turn queries, limited visual modalities, and lack a framework to assess reasoning quality over multiple steps as required in real-world settings. To address this, we introduce Agent-X, a large-scale benchmark for evaluating vision-centric agents’ multi-step and deep reasoning capabilities in real-world, multimodal settings. Agent-X features 828 agentic tasks with authentic visual contexts, including images, multi-image comparisons, videos, and instructional text. These tasks span six major agentic environments: general visual reasoning, web browsing, security and surveillance, autonomous driving, sports, and math reasoning. Our benchmark requires agents to integrate tool use with explicit, stepwise decision-making in these diverse settings. In addition, we propose a fine-grained, step-level evaluation framework that assesses the correctness and logical coherence of each reasoning step and the effectiveness of tool usage throughout the task. Our results reveal that even the best-performing models, including GPT, Gemini, and Qwen families, struggle to solve multi-step vision tasks, achieving less than 50% full-chain success. These findings highlight key bottlenecks in current LMM reasoning and tool-use capabilities and identify future research directions in vision-centric agentic reasoning models. Our data [^1] https://huggingface.co/datasets/Tajamul21/Agent-X and code [^2] https://github.com/mbzuai-oryx/Agent-X are available.
@@ 1 Introduction
Agentic frameworks enable AI systems to interact with their environment by perceiving inputs, invoking tools, and executing actions. While perception and tool use are core components, solving complex tasks effectively requires reasoning, i.e., the ability to draw logical inferences, make decisions, and adapt over time based on multimodal inputs such as text, images, video, and temporal context (Kumar et al., 2025; Yang et al., 2023; Yao et al., 2023; Thawakar et al., 2025). Deep reasoning in agents helps effectively plan, execute, and adapt across evolving scenarios (Liu et al., 2024b; Nathani et al., 2025). Recent works have explored integrating reasoning and tool use within large multimodal models (LMMs)(Achiam et al., 2023; Bai et al., 2023b; Bi et al., 2024; Team et al., 2023; Jiang et al., 2023; Grattafiori et al., 2024), where LMMs act as controllers for planning, and callable tools handle external actions(Chase, 2022; Gravitas, 2023; OpenAI, 2023). This architecture enables agents to combine perception, visual understanding, symbolic reasoning, and generative capabilities, significantly improving task performance in complex environments.
@@
Existing benchmarks for agentic systems have primarily focused on text-based interactions, with limited support for multimodal inputs such as images, videos, and multi-image comparisons (Mialon et al., 2023; Liu et al., 2024b; Wang et al., 2024). While some efforts have extended to multimodal tasks, they are often restricted to static images, synthetic environments, or narrowly scoped domains, offering limited insight into an agent’s ability to perform complex, tool-driven reasoning. A key limitation is inadequate evaluation of deep reasoning [^3] In our context, Deep reasoning refers to coherent, multi-step inference grounded in multimodal context.: current benchmarks either neglect this aspect entirely or lack principled metrics for assessing multi-step logical coherence (Wang et al., 2024; Nathani et al., 2025; Liu et al., 2024b; Team, 2023). Furthermore, most existing benchmarks rely on either fully synthetic tasks or manual annotations, which lack scalability and fail to capture the depth of reasoning or the complexity of video-based multimodal interactions, falling short of modeling real-world agentic tasks (Han et al., 2024; Thawakar et al., 2025), as summarized in Table 1. As agents rapidly evolve in their capabilities, existing benchmarks struggle to keep pace, underscoring the need for more comprehensive, reasoning-centric evaluations to track planning, adaptation, and tool use in authentic, multimodal interactive scenarios (Kiela et al., 2023).
@@
Table 1: Comparison of Agentic Benchmarks. Columns show key dimensions including scale, realism, modality, reasoning depth, tool interaction, and annotation quality. Our benchmark Agent-X uniquely supports all criteria with 828 diverse, manually verified agentic tasks.
@@
Benchmark Agentic Tasks Multimodal Tools Real-world Queries Multimodal Inputs Deep Reasoning Executable Tools Hybrid Annotation
APIBench (Patil et al., 2024)
APIBank (Li et al., 2023)
ToolBench (Qin et al., 2024)
MLAgentBench (Huang et al., 2024) -
GAIA (Mialon et al., 2023) 466
SWE-Bench (Jimenez et al., 2024)
GTA (Wang et al., 2024) 229
MLE-Bench (Chan et al., 2024)
m&m’s (Liu et al., 2024c) -
RE-Bench (Wijk et al., 2024) 7
ScienceAgent (Chen et al., 2024b) 102
MLGym (Nathani et al., 2025) 13
Agent-X (Ours) 828
@@
Agent-X is the first benchmark that combines large-scale, real-world multimodal inputs (images, video, text) with tool-augmented step-wise reasoning evaluation across six environments, providing both breadth and depth missing in GAIA (Mialon et al., 2023), GTA (Wang et al., 2024), and other recent datasets. To address these limitations, we introduce Agent-X, a large-scale evaluation benchmark for vision-centric agents that emphasizes two core principles: multimodal reasoning and vision-first evaluation. Agent-X rigorously tests agents’ ability to process complex visual and textual inputs, execute tool-augmented plans, and perform deep reasoning across real-world tasks. A primary feature of our benchmark is its deep reasoning assessment, i.e., judging the ability to perform coherent, multi-step problem-solving, logical inference, and adaptive planning. To move beyond surface-level evaluation (focused on the correctness of the final result), we introduce fine-grained metrics that capture both intermediate reasoning steps and overall task coherence. This approach helps distinguish genuine logical progression from inconsistencies and confabulations (plausible-sounding but disconnected steps) in reasoning chains, as illustrated in Figure 1. Agent-X spans six diverse multimodal environments, incorporating images, videos, and spatiotemporal contexts to evaluate agents in rich, realistic settings that demand generalization beyond text. Our task pipelines are derived from authentic user queries without explicit step instructions, simulating naturalistic agent interaction as highlighted in Table 2. These tasks are paired with executable toolchains via a semi-automated pipeline that integrates a broad set of real-world tools, enabling scalable tool-augmented decision-making. In Agent-X, these components create a comprehensive framework for evaluating agentic performance in perception, reasoning, and action. Our main contributions are as follows:
@@
ToolBench I’m writing a blog post… first retrieve available figlet styles and then generate ASCII art for the strings using the style. figlet, list figlet styles
APIBench I am an engineer at Uber and… Write a python program in 1 to 2 lines to call API in TorchHub. ObjectDetection
GTA How much should I pay for the soda in the picture according to the price on the menu? ImageDescription, CountGivenObject, OCR
m&m’s I need an illustration for my children’s book. I’ve imagined a scene where… After we have the image, we also need to identify all the objects, then add labels to them. ImageGeneration, ObjectDetection, Tagging
@@ 2 Related Work
Large Multimodal Agents: The rapid progress in LMMs has catalyzed the development of agentic frameworks capable of autonomous planning, tool use, and decision-making. A growing body of research has explored the extension of LMMs beyond conventional text generation by integrating them with external tools such as APIs (Zhang et al., 2025b), document processing (Musumeci et al., 2024), operating systems (Mei et al., 2024), and web interfaces (Song et al., 2024), allowing them to interact meaningfully with their environments. This evolution has led to the emergence of tool-augmented agents such as Avatar (Wu et al., 2024), LangChain (Chase, 2022), AutoGPT (Gravitas, 2023), and BabyAGI (Nakajima, 2023), which provide frameworks for textual reasoning and execution of external actions. In the context of web browsing, systems such as WebShop (Yao et al., 2022), WebGPT (Nakano et al., 2021), and WebCPM (Qin et al., 2023) enhance LMMs with capabilities for browsing, searching, and information retrieval. Beyond textual modalities, integrations such as RestGPT (Song et al., 2023) and AppAgent (Zhang et al., 2025c) enable interaction with REST APIs and emulate touchscreen operations, respectively. In the multimodal domains, vision-centric agents like MLLMTool (Wang et al., 2025) and LLaVA (Liu et al., 2024a) equip LMMs with the ability to reason about visual content through pre-trained vision models. Nevertheless, there is no standardized evaluation protocol to rigorously assess the reasoning capabilities of these agents, particularly in tasks that require tool invocation and decision-making.
@@
Multimodal Agentic Benchmarks: A broad range of recent benchmarks have emerged to evaluate the performance of agentic frameworks. Recent efforts like ToolBench (Qin et al., 2024) introduce REST APIs and evaluate tool use with metrics such as pass rate and win rate, while APIBench (Patil et al., 2024) assesses accuracy for APIs. Complementing these, APIBank (Li et al., 2023) presents a diverse suite of commonly used APIs, such as search engines and hotel reservations, offering a comprehensive framework for evaluating the planning, retrieval, and execution skills of agents based on LMM. Benchmarks such as ToolQA (Zhuang et al., 2023), Gentopia (Xu et al., 2023), Gorilla (Patil et al., 2024), and AgentBench (Liu et al., 2024b) often repurpose standard data sets or follow restricted evaluation protocols, which limit their utility in assessing open-ended and generalizable tool use capabilities. To address this, GAIA (Mialon et al., 2023) poses conceptually challenging questions for human-like understanding, GTA (Wang et al., 2024) evaluates agents with executable toolchains in real-world settings, and OSWorld (Xie et al., 2024) offers multi-step tasks based on real user behavior. MLGym (Nathani et al., 2025) further adds complex multimodal tasks requiring deep reasoning and tool use. As shown in Table 1, despite promising advances, current evaluations face two key limitations: (1) most benchmarks focus on final answer accuracy, overlooking interpretability, which is crucial for agentic tasks involving reasoning and tool use, and (2) the absence of standardized and interpretable evaluation protocols hinders meaningful comparison across methods.
@@
Reasoning in Large Multimodal Models: Multiple methods have been introduced to enhance the reasoning capabilities of LMMs, enabling models to handle complex, multi-step problems through explicit reasoning processes (Amizadeh et al., 2020). Following the rise of reasoning techniques, benchmarks have emerged to evaluate reasoning and chain-of-thought in models. Early efforts like CLEVR (Johnson et al., 2017) tested basic visual logic, while StrategyQA (Geva et al., 2021) introduced multi-hop reasoning. ScienceQA (Saikh et al., 2022) added multimodal scientific tasks, MathVista (Lu et al., 2023) unified math reasoning datasets, and ShareGPT-4o (Zhang et al., 2024) offered extensive chain-of-thought samples to enhance intermediate reasoning. The m&m’s (Liu et al., 2024c) benchmark focuses on multi-step, multimodal reasoning and evaluates various planning strategies. However, many benchmarks, including m&m’s, rely on AI-generated queries with predefined tool sequences, limiting their realism. LLaMA-V (Thawakar et al., 2025) introduced a benchmark focused on assessing tasks requiring multiple reasoning steps. While these reasoning-focused benchmarks have been introduced, they do not address the agentic framework or evaluate tool usage and the associated reasoning within vision-centric agentic tasks.
@@ 3 Agent-X Benchmark
The Agent-X benchmark evaluates the capacity of vision-centric agents to perform complex reasoning tasks that necessitate proficient tool usage across diverse, real-world scenarios. It accommodates a broad spectrum of input modalities, including single-image analysis, multi-image comparisons, video interpretation, and multimodal image-text interactions. Agent-X presents agents with authentic visual data and requires them to use multi-step reasoning processes. The construction of Agent-X follows a semi-automated pipeline (illustrated in Figure 2), where an initial set of candidate queries is generated through large multimodal models (LMM) and subsequently refined and validated by human experts.
@@ 3.1 Benchmark Design
In our framework, $\mathcal{T}_{c}$ comprises a carefully curated collection of 14 tools distributed across diverse environments. These 14 tools capture the breadth of capabilities required for real-world, vision-centric reasoning. A comprehensive list of these tools is presented in Appendix §B. The queries $\mathcal{Q}$ are classified into three distinct categories: factual, interpretive, and generative. Illustrative examples of these query types are provided in Appendix §I. In the case of a factual query, the answer is a specific, uniquely determined value, such as a number or a phrase. For an interpretive query, the final answer consists of descriptive text. While the answer is not unique, it conveys a general concept or idea, with the final response comprising a reference answer. For a generative query, we do not directly utilize the generated output since the LMMs are descriptive models; they only provide textual descriptions rather than generating visual content like images/videos. In these instances, $\mathcal{A}_{i}=\emptyset$.
@@ 3.2 Task Pipeline
Our query construction process begins with generating candidate queries using an LMM, which is provided with the visual input $\mathcal{V}_{i}$ and the available complete toolset $\mathcal{T}_{c}$. The LMM generates initial queries that human annotators then refine and validate for clarity, coherence, and realism. In the next step, LMM is re-engaged with the refined query to produce a detailed reasoning trace $\mathcal{R}_{i}$, capturing the sequence of tool calls (input/output), intermediate decisions, and final answer $\mathcal{A}$. Human reviewers further ensure that the trace is logically consistent and aligned with the query, while the final answer, $\mathcal{A}_{i}$ and its justification, $\mathcal{J}_{i}$ are validated for correctness and comprehensibility. The prompts provided to the LMM for query generation and reasoning trace construction are included in Appendix §J.